prettyariannaher11:

can u solve it for me x divide by 5 equal 2

CripQUEZZ:

it would be 10? its simple math multiply 5 by the 2 u get x

prettyariannaher11:

thank u i got right

xxDeppressionxx:

@cripquezz wrote:
it would be 10? its simple math multiply 5 by the 2 u get x
Nice work!, 👍

xxDeppressionxx:

Lol even though it was easy

itsmehjay:

An easy way to find the answer is to do the opposite operation which would be multiplying but yeah good job

prettyariannaher11:

what is 144=12z

CripQUEZZ:

@prettyariannaher11 wrote:
what is 144=12z
divide the 144 by 12 u jus do the opposite opperation so 144/12 = ?

prettyariannaher11:

i ask what the answer is

CripQUEZZ:

@prettyariannaher11 wrote:
i ask what the answer is
ok i gotchu 12 x 12 = 144 so u jus do the opposite 144/12 = 12

CripQUEZZ:

12 = z

prettyariannaher11:

thanks

prettyariannaher11:

what is the answer y-4.6=12

xxDeppressionxx:

Please close this question, if you are asking question about simpler problems you may want to consider using an online calculator. Also when you need help with these and you can’t solve them ask an adult around you.

prettyariannaher11:

stop being mean

CripQUEZZ:

ain mean jus the truth

prettyariannaher11:

then you dont have to answer

xxDeppressionxx:

QuestionCove does not allow problems like, 1+1 or 100*100 those are simple to answer just use a multiplication chart or calculator.